### v4

Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to capture time-varying information, particularly for widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose a fresh approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation and measurements for OV tasks. Building upon the prior alignment(s) from OM, we propose a pipeline optimisation method called the cross-reference (CR) mechanism to enhance overall OV performance. We experimentally validate the OM4OV pipeline and the cross-reference mechanism in an OV testbed originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets. We also discuss insights into OM used for OV tasks, where some apparent false mappings detected by OV systems are not actually untrue.
keywords:
ontology matching, ontology versioning
category:
[…]
1 Introduction
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [8]. An ontology is not static, and the need for version control arises with its origination. While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [13]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [9]. This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
Ontology versioning (OV) aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology can apply the correct changes correspondingly [13]. While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new annotation to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of free-text notebooks, an extensible markup language (XML) document, or a knowledge graph (KG). However, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capabilities in detecting incorrect version information or identifying missing version information.
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to address interoperability between ontologies, with ontology entities serving as inputs. While OM is a well-studied problem [17], a unified pipeline that can extend and reuse existing OM techniques and systems for OV tasks would be highly desirable. However, OM techniques and systems cannot be directly applied to OV tasks because OM and OV have some significant differences. The OM input is two distinct ontologies, while the OV input is expected to be two different versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises two sets: changed entities and unchanged entities. OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions of one ontology. To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks, thereby leveraging the deep research legacy of OM to assist with OV. Specifically, our key contributions include:
  • We systematically analyse OV from an OM perspective and introduce a novel OM4OV pipeline.
  • Drawing on OM practice, a novel cross-reference mechanism is proposed to optimise candidate selection and improve the overall performance of the OM4OV pipeline.
  • We implement the OM4OV pipeline and the cross-reference mechanism in a proof-of-concept system and experimentally evaluate its performance.
  • We argue that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of a similarity threshold.
The remainder of the paper is organised as follows. Section 2 reviews the literature on OV. Section 3 introduces our novel OM4OV pipeline, while Section 4 evaluates the proposed pipeline. We propose the cross-reference mechanism as a pipeline optimisation method and experimentally evaluate its performance in Section 5. Section 6 and 7 discuss our implications and current limitations. Section 8 concludes the paper.
2 Related Work
Version control is recognised as a vital element in ontology management. Different versions of an ontology need to be interoperable so that version changes do not impede the effective and sustainable use of the ontology. There have been two main approaches towards OV that aim to enhance an ontology with the ability to represent different versions and to identify their differences.
One option is to include version information within the ontology. The Simple HTML Ontology Extensions (SHOE) [11] uses the tag BACKWARD-COMPATIBLE-WITH to record version information. The authors in [12] argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language [4], where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify an entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare deprecated entities. Later, the ontology language $\tau$OWL [22] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [20].
A second option is to create a separate version log to track changes in versions. Unlike traditional approaches that use an unstructured plain text file, the authors in [18] propose a new approach that uses a version ontology with the change definition language (CDL) to create a version log. In [5], the authors construct a historical knowledge graph (HKG). Storing the version log in the knowledge graph not only avoids repetition, but also enables advanced search functions. The authors in [21] argue that the version logs may contain redundancy and inconsistent information. They propose a graph-of-relevance approach for interlinking different version logs and removing less relevant versions.
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. To the best of our knowledge, our work is the first to systematically analyse and utilise OM for OV. The output of our novel automatic OV detection could be easily recorded by any of the current approaches presented above.
3 OM4OV Pipeline
Figure 1 illustrates the overview of the OM4OV pipeline. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), an OM task can be considered as finding an alignment ($A$) that contains a collection of mappings. Similarly, given an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), an OV task can be considered as finding an alignment (A) that contains a collection of mappings over the two different versions. However, while an alignment for OM only considers matched entities, OV tracks both matched and non-matched entities. Furthermore, matched entities are composed of two subsets remain and update, while non-matched entities are composed of add and delete entities. In practice, we expect that unchanged remain entities dominate for OV.
Figure 1: Overview of the OM4OV pipeline.
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as [6]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$. Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. While OM may have different mapping relations, OV narrows down the task and focuses only on the equivalence relation.
Definition 3.1.
An OV task is defined as:
$A_{match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},s\leq c\leq 1\}$ (1)
$A_{non-match}$ =$=\{(e1,e2,\equiv,c)|e1\in O,e2\in O^{\prime},0\leq c<s\}$
We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$). The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
Definition 3.1.1.
$A_{match}$, $A\odot$ and $A\otimes$ are defined as:
$A_{match}=A\odot\cup A\otimes$ (2)
$A\odot=A_{match}(c=1)$
$A\otimes=A_{match}(s\leq c<1)$
The add and delete entities are actually non-matched entities between different ontology versions. The add entities are the non-matched terms in $O^{\prime}$, which have no matches in $O$. Similarly, we can interpret the delete entities as non-matched entities in $O$.
Definition 3.1.2.
$A_{non-match}$, $A\oplus$, and $A\ominus$ are defined as:
$A_{non-match}=A\oplus\cup A\ominus$ (3)
$A\oplus=\{e2\in O^{\prime}\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$
$A\ominus=\{e1\in O\phantom{{}^{\prime}}\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision (Prec), recall (Rec), and the F1 score (F1). While precision measures matching correctness and recall measures matching completeness, the F1 score offers a harmonic mean to balance correctness and completeness. OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance.
Definition 3.2.
The measures of OV task are defined as:
$Prec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot|\otimes|\oplus|\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
4 Pipeline Evaluation
4.1 Dataset Construction
There is a dearth of benchmark datasets for the evaluation of OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Given a source ontology and a target ontology with their reference alignment reference.xml, our approach is described in the following steps:
Figure 2: Generation of OM4OV datasets.
{bracketenumerate}
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned to one of these. For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have the four sets as pairwise disjoint.
For entities assigned to update, we need to generate a new name for the updated entity. We should expect the new name to have a similar meaning to its original name. For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”. To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus. For entities whose names are unique identifiers or codes (and not textually meaningful names), we use their annotation properties (e.g. rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead of the code. For entities in update that do not have synonyms in the corpus, we randomly re-assign them to remain, add, or delete.
Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml. For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list. For entities assigned to remain, no operation is required.
We generate $O$ and $O^{\prime}$ according to the following rules: {alphaenumerate}
$O$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{add or }p\in\textit{add or }o\in\textit{add}\}$. That is, $O$ is constructed as $O_{i}$ without all the triples related to entities in the add list.
Let $e$ be an entity and $M$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$. Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in M$ and to be $e$ otherwise. Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$. That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities. (2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities. Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. For reproducibility, we use a fixed seed to control the randomness in steps (3) and (4).
4.2 Experiment Setup
Table 1 lists the detailed information of the selected OAEI track used for the OV testbed. The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies. Figure 3 shows the number of entities in each track. The fixed seed for the OV testbed is set to 42.
Table 1: Selected OAEI tracks for the OV testbed.
Track Domain Alignment Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
Figure 3: Number of entities.
The OM4OV pipeline is implemented in Agent-OV, as an extension of Agent-OM [19]. Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [16] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [14], Alibaba Qwen [1], Google Gemma [7], and ChatGLM [23]. For performance using different LLMs, we refer the reader to [19], where we find that API-accessed LLMs generally perform better than open-source LLMs. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV.
4.3 Results
Figure 4 evaluates Agent-OV and Agent-OM on the OV testbed. The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated. The results indicate that it is possible to unify the OM and OV tasks as developed in this paper. Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 scores of Figure 4. However, the original OM systems require necessary modifications to tackle OV tasks. We can see that Agent-OM always has an unusually high overall performance across all alignments. This is because the traditional OM system only captures matched entities between two ontologies, where this refers to the remain and update entities in the OV task. While remain entities commonly comprise the major proportion of entities over two different versions, the measure of remain entities dominates the overall evaluation. We believe this is a common situation and causes misleading performance measurement when using the OM system in OV tasks. Agent-OV overcomes the skewed measure in Agent-OM. When decomposing the measure into four sub-measures, we can observe a more precise matching performance across different OV tasks. In general, we observe the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. {bracketenumerate}
The measures for remain are typically very close to 100%. This sub-measure is statistically significant and can be omitted in OV tasks.
The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities.
The measures for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Additionally, we observe a longer computation time in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks. There is a need to optimise the pipeline for OV in large-scale ontologies.
Figure 4: Evaluation of Agent-OM and Agent-OV on the OV testbed.
Figure 5 studies the effect of varying similarity thresholds in Agent-OV. We change the similarity threshold from 0.9 to 1 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similarly to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Furthermore, detecting update entities and detecting add and delete operations are negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities. Our explanations for these trends are as follows.
Figure 5: Evaluation varying similarity thresholds in Agent-OV. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
{bracketenumerate}
Within each sub-measure, the precision-recall trade-off still holds. Across different sub-measures, they are not independent but satisfy the following equations, where $N(O)$ is the number of entities in $O$ and $N(O^{\prime})$ is the number of entities in $O^{\prime}$:
$N(O)+N(O^{\prime})$ $=2\times(|A\odot|+|A\otimes|)+|A\oplus|+|A\ominus|$ (5)
For each change in a part of an alignment, other parts will change accordingly. For example, if a new alignment is found in $A\odot$ or $A\otimes$, then the number of alignments in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa. If we define $\Delta A$ as a universal change in OV, any changes in the alignments satisfy the following equation:
$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=|\Delta A\oplus|=|\Delta A\ominus|$ (6)
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$. For recall, there are direct relations between sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa. For precision, there are no direct relations between sub-measures. The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or uncertain ($\sim$), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
$Rec\odot|\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|R|}\quad Prec\odot|\otimes\sim=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
$Rec\oplus|\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|R|}\quad Prec\oplus|\ominus\sim=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
5 Pipeline Optimisation
5.1 Motivation
Often, ontology creators provide cross-references to other ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the CMT ontology and the Conference ontology is provided alongside the CMT ontology. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
5.2 Approach
Figure 6 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be incorporated into the final alignment, while the known non-matching regions will be removed entirely in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also enhanced by utilising these established mappings. Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this simplifies the detection of the posterior alignment.
Figure 6: Cross-reference mechanism.
Definition 5.1.
Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,s\leq c\leq 1\}$ (8)
$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},s\leq c\leq 1\}$
Definition 5.2.
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{\pi}$ and $A_{\pi^{*}}$.
$A_{match}=A_{\pi}\cup A_{\pi^{*}}$ (9)
Definition 5.2.3.
$A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (10)
$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
Definition 5.2.4.
[…]
$A_{\pi}$ $=(R_{or1}\cap R_{o^{\prime}r1})\cup(R_{or2}\cap R_{o^{\prime}r2})\cup...\cup(R_{orn}\cap R_{o^{\prime}rn})$ (11)
$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
5.3 Evaluation
Figure 7 compares the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
Figure 7: Precision, Recall and F1 evaluation of the cross-reference (CR) mechanism on the OV testbed.
Figure 8 compares the effects of different similarity thresholds in the CMT Ontology versioning. The results of the experiment show that our cross-reference mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline, which means the cross-reference mechanism can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
6 Discussion
So far, we have experimentally validated the OM4OV pipeline with a novel cross-reference mechanism for pipeline optimisation. Although matching performance has been improved, it still does not achieve 100% as false mappings may still exist. Are these mappings actually false?
{bracketenumerate}
False mappings in OV can result from a flawed ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by inconsistent naming convention practice. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
cmt:hasAuthor => cmt:writtenBy
cmt:writtenBy => cmt:isWrittenBy
False mappings in OV can create an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in domain understanding or evolution of natural language in the domain.
cmt:Chairman => cmt:Chair
cmt:ConferenceChair => cmt:General_Chair
False mappings in OV can arise from an inappropriate setting of the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure each term is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
s = 0.95, cmt:SubjectArea => None
s = 0.90, cmt:SubjectArea => cmt:Topic
7 Limitations
We focus mainly on tracking conceptual changes in classes and properties (e.g. adding, deleting, or updating class and properties). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.
While OM has a universal measure, our proposed measures for OV have four sub-measures. It is necessary to find a single universal measure combining these sub-measures so that we can fairly assess and compare the system performance in OV and OM. We plan to investigate a merged formula for OV sub-measures in the future. For example, using a harmonic mean to combine sub-measures.
8 Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with minor modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks. Evaluation experiments on our OAEI testbed validate the feasibility and reliability of our system. Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods, that is, using URIs or additional versioning triples, and ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, smart building ontologies Brick Schema [3], RealEstateCore [10], and ASHRAE Standard 223P [15].